### v2

# “Feeling that I was Collaborating with Them:” A 20-year Systematic Literature Review of Social Virtual Reality Leveraging Collaboration

Abstract.
As more people meet, interact, and socialize online, Social Virtual Reality (VR) emerges as a promising technology that can bridge the gap between traditional face-to-face and online communication. Compared to traditional screen-based applications, Social VR provides immersive, spatial, and three-dimensional social interactions, making it a promising tool for enhancing collaborations. To map the existing research in this domain, we conducted a 20-year systematic literature review to characterize how Social VR has been employed for collaboration. After screening 2,035 articles, we identified 62 articles that addressed how Social VR has supported collaboration among remote users. Our findings show that Social VR can enhance team collaboration on three key levels: enhancing individual perceptions and experiences within their groups, fostering team dynamics with virtual elements that enable realistic interactions, and employing affordances unique in VR that augment users’ spaces. Future research should explore how Social VR can support long-term collaboration, foster trust, enable more diverse and inclusive participation, and move beyond replicating physical-world interactions by leveraging the unique affordances of immersive environments. This review highlights the current practices, challenges, and future research opportunities within CSCW, offering insights for theorizing the impact of Social VR on team collaboration and for designing new applications that effectively support remote collaborations.
Social VR, Virtual Reality, Collaboration, Extended Reality, Human-Computer Interaction
copyright: acmlicensedjournal: PACMHCIjournalyear: 2025journalvolume: 4journalnumber: CSCW2publicationmonth: 10price: 15.00ccs: Human-centered computing Collaborative and social computingccs: Human-centered computing Human computer interaction (HCI)
[…]
Figure 1. Overview of the systematic literature review findings on Social Virtual Reality (VR) collaboration. The figure illustrates the three main themes: (1) Individual Perception and Experience of Collaboration in Social VR, examining sense of presence and avatar perception; (2) Team Dynamics and Collaboration in Virtual Environments, focusing on social presence, collaboration realness, trust-building, and engagement; and (3) Enabling Collaboration through Social VR Affordances, addressing virtual environment functionality, virtual objects and tool facilitation, and immersion enhancement through spatial audio and movement.
1. Introduction
The rise of Social Virtual Reality (Social VR) is reshaping how people can communicate and collaborate in digital spaces. In February 2022, Meta’s social VR Horizon Worlds achieved 300,000 active monthly users in the US (Heath, 2022) and the platform Rec Room surpassed 3 million monthly active users (Feltham, 2022), showing the growing adoption of Social VR as an environment for social connections.
Unlike traditional screen-based communication tools, such as video conferencing or instant messaging, Social VR enables users to immerse themselves in three-dimensional spaces through head-mounted displays (HMDs), creating a sense of physical presence and embodied interaction in simulated environments (Maloney and Freeman, 2020). Social VR offers unique affordances that can enhance communication among distributed groups, such as avatar embodiment, shared object manipulation, spatial audio, and interactive environments (Li et al., 2021; Freeman et al., 2022). Its capacity to simulate real-world experiences has positioned Social VR as a transformative medium for social interactions, particularly after the COVID-19 pandemic, which increased the need for robust alternatives to face-to-face interactions (Waizenegger et al., 2020; Ball, 2022; Gómez-Zará et al., 2023).
Despite the growing interest in Social VR, its use in collaborative and task-oriented settings remains unclear. While much of the existing research has focused on its applications in gaming, training, and casual gatherings (Bujić et al., 2021; Angel-Urdinola et al., 2021; Maloney and Freeman, 2020), the potential of Social VR for enabling collaborative meetings and teamwork has been underexplored. Companies such as Microsoft have attempted to integrate Social VR environments into professional settings to support meetings, conversations, and teamwork (Tilley, 2022; Roth, 2023). However, adoption has often fallen short due to several challenges, including low-quality simulated environments, steep learning curves, cognitive dissonance between physical and virtual worlds, and a lack of applications specifically designed for professional purposes (Ashtari et al., 2020; Zhao, 2011; Hubbard and Aguinis, 2023). Moreover, the literature has yet to fully explore how these immersive technologies support team processes, such as trust-building, long-term collaboration, and group cohesion (Maloney and Freeman, 2020; Abramczuk et al., 2023; Osborne et al., 2023). These challenges highlight important gaps in both the design and understanding of Social VR as an environment for distributed collaboration. Specifically, there is a need to better characterize how Social VR environments can address these limitations and replicate the richness of in-person interactions in group settings.
Characterizing studies on Social VR that address collaboration and teamwork is important to understanding its potential as a transformative tool for distributed collaboration. Research communities in Computer-Supported Cooperative Work (CSCW) and Human-Computer Interaction (HCI) have long sought to design and study systems that support collaboration among remote groups (Olson and Olson, 2000; Harris et al., 2019). As such, Social VR offers an opportunity to rethink how digital tools can enhance collaboration in remote and hybrid contexts, bridging the limitations of current systems while using its unique affordances for presence and engagement. This paper provides a comprehensive overview of the current research landscape in Social VR, identifying critical gaps in areas such as trust-building, user engagement, scalability of teamwork, long-term collaboration, the diversity of study samples, and the design of effective applications to support various stages of collaborative processes. By synthesizing these findings, we aim to highlight actionable opportunities for future CSCW research and offer design implications for developing effective collaborative VR systems that address the limitations of existing tools. This effort is intended to bridge the gap between theoretical advancements and practical applications, helping organizations and researchers leverage Social VR for more meaningful and productive remote teamwork. Given this motivation, the research questions that guide this study are:
  • RQ1: How have previous studies addressed team collaboration in Social VR?
  • RQ2: What are the specific gaps, future prospects, and research directions in this area?
To address these research questions, we conducted a systematic literature review on Social VR addressing aspects of collaboration. Our goal is to identify key solutions, insights, and recommendations that support group interaction and teamwork in Social VR contexts. This systematic review covered articles published in the IEEE Xplore Digital Library, ACM Digital Library, and Web of Science databases from 2004 to 2024, covering 20 years of research. After screening 2,035 articles, we identified 62 articles that specifically focused on how Social VR has supported teams and enhanced collaboration. Based on this corpus, we conducted a thematic analysis and identified three core themes: (1) “Individual Perception and Experience of Collaboration in Social VR,” (2) “Team Dynamics and Collaboration in Virtual Environments,” and (3) “Enabling Collaboration through Social VR Affordances.” These themes illustrate how Social VR’s affordances and mechanisms at the individual, team, and environment levels can facilitate collaboration.
Our systematic review also highlights several research gaps, including insufficient attention to diverse user populations across demographics and technical expertise, limited understanding of how Social VR scales to larger teams and complex tasks, and a lack of standardized evaluation frameworks for assessing collaboration. Moreover, the themes uncover key questions about trust-building mechanisms, team dynamics, and the transition from traditional online platforms to Social VR for collaboration purposes. Overall, this review provides an comprehensive assessment of the current capabilities and promising directions that Social VR can offer to people and organizations in real-world collaborative settings.
The contributions of this paper are twofold. First, it offers a comprehensive review of the use of Social VR environments to support remote collaboration, encompassing an analysis of applications and their impact on teamwork. Second, it provides a holistic map for future CSCW research and system development in Social VR to enhance collaboration. The themes and implications discussed in this study aim to inform CSCW designers and researchers about the opportunities and challenges involved in making Social VR effective for remote collaboration.
2. Background
We situate our work around prior studies examining (a) teams and technologies to support teamwork, (b) Social VR, and (c) how people have used Social VR to collaborate in task-oriented settings.
2.1. Teams and Teamwork Technologies
Teams are groups of two or more individuals working interdependently, adaptively, and dynamically toward a common and valued goal, bringing diverse skills, knowledge, and perspectives to address challenging tasks (Gómez-Zará et al., 2020; Salas et al., 2000). Unlike individuals working alone, teams can bring together people with diverse skills and knowledge to achieve better outcomes (Engel et al., 2015; Mathieu et al., 2017). Team members work on different parts of a task, relying on each other, sharing responsibility to achieve a common goal, and resolving complex problems in a distributed manner (Mathieu et al., 2017; Gómez-Zará et al., 2020; Salas et al., 2000). Previous research shows that successful teamwork relies on clear communication, coordination, mutual support, and individual contributions aligned with the overall task strategy (Mathieu et al., 2017). While teamwork requires more coordination and collaboration than individual efforts, it provides a higher capacity to manage challenging tasks, particularly when specialized skills are needed (Mathieu et al., 2017). The importance of teams has grown as organizational structures have shifted over the past decades from individual roles to team-based systems (Wuchty et al., 2007), with tasks increasingly designed around groups working toward shared objectives (Arora et al., 2023; Mathieu et al., 2017). For instance, in IT organizations, project teams require programmers, designers, and project managers to work interdependently on complex software development projects, where effective communication, task coordination, and shared understanding are essential for delivering successful outcomes (Arora et al., 2023; Huang et al., 2023).
The impact of online technologies on collaboration and teamwork continues to evolve alongside the dynamic technological ecosystem (Dulebohn and Hoch, 2017; Jimenez et al., 2017; Treem et al., 2020). Advancements in groupware, intranets, social platforms, smartphones, cloud technologies, and videoconferencing platforms have redefined the degrees of “virtualness” (Staples and Webster, 2008; Brown et al., 2020) in which teams operate, ranging from purely face-to-face to only virtual interactions (Griffith and Neale, 2001; Brown et al., 2020). For example, hybrid workplaces allow team members to work remotely and meet in person on specific occasions (Brown et al., 2020). Workers can still be physically in one building, but coordinate through online communication. Since team members do not necessarily work in the same physical place, current challenges that teams face reside in coordinating, staying cohesive, and performing effectively (Gilson et al., 2015).
With the rise of hybrid and remote work (Freeman et al., 2022; Bjørn et al., 2014; Breideband et al., 2023), online collaboration tools are now key components for facilitating effective communication and teamwork (Cai and Tanaka, 2019; Li et al., 2021). Instant messaging platforms, like Microsoft Teams and Slack, provide another layer of group communication by enabling quick exchanges of information and facilitating asynchronous collaboration (Cai and Tanaka, 2019; Li et al., 2021). These online communication platforms support various forms of media sharing and integrate with other productivity tools, enhancing workflow efficiency and group awareness. However, they may contribute to information overload and can sometimes lead to misinterpretations without the contextual cues present in verbal communication (Cha et al., 2018). Video conferencing platforms (e.g., Zoom, Google Meets, Microsoft Teams) also create a shared visual space that enhances user engagement and presence, facilitating real-time communication from different locations (Tang et al., 2023; Karis et al., 2016). Despite offering significant benefits, they often lack the depth of non-verbal cues available in face-to-face interactions, such as body language (Freeman and Acena, 2021; Bente et al., 2008; Kanawattanachai and Yoo, 2002), and can increase the mental exhaustion and cognitive load for their prolonged use (Bailenson, 2021). Previous research has also shown that remote collaboration can curb several activities that require group thinking, such as creativity (Brucks and Levav, 2022), idea generation (Lin et al., 2023), identity (Bartel et al., 2007), and cohesion (Yang et al., 2022). These current limitations have motivated CSCW researchers to explore how immersive environments, such as Social VR, can improve remote collaboration and enhance the quality of remote teamwork.
2.2. Social VR
Social Virtual Reality (VR) refers “to 3D virtual spaces where multiple users can interact with each other using VR head-mounted displays (HMDs)” (Maloney and Freeman, 2020). Social VR encompasses a subset of VR platforms and applications specifically designed to support social interaction and co-presence (Li et al., 2021; Freeman and Maloney, 2021), where users can feel as if they are physically present with others in the same virtual space (McVeigh-Schultz et al., 2019). Given these unique affordances, Social VR has the potential to change how people engage, communicate, and collaborate in various activities, from social gatherings and entertainment to educational and professional teamwork settings (Deighan et al., 2023). Platforms like ‘VRChat’ have become popular spaces for individuals to connect in VR environments and foster a sense of community and well-being, particularly with the emergence of the COVID-19 pandemic (Osborne et al., 2023).
Unlike traditional video conferencing applications, Social VR allows people to immerse themselves in simulated environments to interact and socialize in more interactive ways (Freeman and Acena, 2021). Social VR applications can enable remote users to interact with 3D virtual objects in immersive spaces (Freeman et al., 2022; Li et al., 2021), utilize avatars that can differ from their real bodies and appearances, employ interactive tools that break the physical laws (Freeman et al., 2022), and help people stay more engaged in the team conversation using their bodies and movements (Abramczuk et al., 2023; Olaosebikan et al., 2022; Steinicke et al., 2020). The unique affordances that VR provides can help users communicate not only with words but also with gestures, body language, and eye gaze with the help of body tracking technologies (Freeman et al., 2022).
Previous research shows that Social VR can enable people to interact in similar ways to face-to-face settings, helping people build stronger social bonds (Yoon and Leem, 2021). For entertainment purposes, VR has demonstrated its potential to enhance shared experiences. For example, Lee et al. (Lee et al., 2022) developed a VR lobby for digital opera social experiences, highlighting how Social VR can replicate and enrich live entertainment through immersive environments. Beyond entertainment, Young et al. (Young et al., 2023) demonstrated that Social VR can facilitate remote learning by enhancing students’ participation in online discussions and activities, improving their sense of presence that traditional online learning platforms often do not provide. In another study, Sanaei et al. (Sanaei et al., 2023) conducted a controlled experiment to compare groups working on a circuit repair in face-to-face (F2F), video conferencing, and VR. The study found that participants in VR had a significantly better task experience and lower cognitive efforts compared to participants in the video conferencing condition. These studies suggest that VR can emulate the interactive aspects of in-person interactions that traditional online communication platforms lack.
Although Social VR has seen increasing use across various settings, its application for remote work and as an online communication tool for group collaboration has not been fully examined.
2.3. Collaboration Work in Social VR
Collaboration in Social VR refers to two or more individuals working together within a Social VR application to achieve a common goal (Freeman et al., 2022). The distinct affordances provided by Social VR, such as avatar customization, nuanced non-verbal behaviors, and spatial interactions, can facilitate a strong sense of being together and foster relationship building among distributed collaborators (Freeman and Acena, 2021). It also enables remote teams to engage in a wide range of collaborative activities that mirror in-person teamwork, from creative projects and problem-solving to mundane tasks and social experiences (Freeman et al., 2022). Furthermore, Social VR allows the interplay of productivity and sociality that often characterizes co-located collaboration (Maloney and Freeman, 2020). In one study, Maloney et al. (Maloney and Freeman, 2020) conducted interviews to explore what makes activities on Social VR platforms meaningful to their users. They found highlighted immersive interactions through full-body tracking and facilitating more natural and expressive communication as the main advantages of Social VR, overcoming the limitations of 2D video conferencing. In another study, Li et al. (Li et al., 2021) argue that Social VR affords more natural social interactions than video conferencing, such as the ability for users to organically break off into small groups and interact with virtual objects in the shared environment. Unlike video conferencing, which restricts users to ’talking head experiences’—static, front-facing video feeds—Social VR allows for more physical activities and spontaneous collaborations that naturally arise from users’ interactions (Li et al., 2021).
Since the COVID-19 pandemic accelerated the shift to remote collaboration, many conference organizers transitioned to online events with mixed success in replicating the traditional experiences, exposing both the strengths and weaknesses of existing conferencing tools (Moss et al., 2021). Several technology companies and research communities explored metaverse technologies—including Social VR platforms—to support synchronous interactions and socialization within 3D virtual environments (Cheng et al., 2022; Duan et al., 2021; Gómez-Zará et al., 2023). Platforms like “Altspace VR” have demonstrated their ability to foster social presence, enabling remote participants to experience a sense of togetherness within a shared, simulated space (Pizzolante et al., 2023). Yet, researchers have also pointed out potential challenges, including accessibility issues, privacy concerns, and the need to establish new social norms for virtual interactions occurring in immersive spaces (Moss et al., 2021; Gómez-Zará et al., 2023).
Given the growing body of research on Social VR, we aim to systematically map how previous research has studied these technologies in the context of collaborations and teamwork. As the CSCW community has delved into several questions and reflections on Social VR, synthesizing key findings and implications from these studies can bring new opportunities for understanding the specific benefits of Social VR for team dynamics, designing more effective collaboration applications, and understanding the barriers to adoption. As technologies continue to shape how people work together (Harris et al., 2019), the CSCW community is well positioned to examine how Social VR’s unique capabilities and affordances enable remote collaborations and guide the development of future technological ecosystems (Wallace et al., 2017; Gonzales et al., 2015).
3. Methodology
We followed a scoping review methodology (Arksey and O’Malley, 2005) to synthesize existing knowledge and reveal trends in using Social VR for collaboration purposes. This methodology allows researchers to map the literature and identify gaps in a specific area of research (Arksey and O’Malley, 2005; Levac et al., 2010). Scoping reviews have several advantages: they are transparent, comprehensive, less prone to bias, and make it easier to reproduce the detailed information reported about each step of the review and how it is conducted (Pham et al., 2014). They have gained prominence in the HCI and CSCW communities as they allow for the synthesis and comprehension of the development of different research directions (Wallace et al., 2017; Cosio et al., 2023; Bergram et al., 2022; Rogers et al., 2022; Wei et al., 2022). We used the Preferred Reporting Items for Systematic Reviews and Meta-Analyses (PRISMA) guidelines to report the methods and results (Page et al., 2021). PRISMA covers four stages in the review process: Identification, Screening, Eligibility, and Inclusion. We collected papers and tabulated them in a shared Google Spreadsheet, capturing metadata such as publication year, abstract, and keywords. Once the metadata was recorded, the authors of the articles were hidden to avoid potential bias during the coding phase.
3.1. Search Strategy and Data Sources
To identify relevant articles for our systematic literature review, we developed a comprehensive search query that could capture the key concepts of our study while balancing sensitivity and specificity. After several iterations, we built our search query by breaking down our research topic into its core components: the technology (i.e., social VR), the activity (i.e., collaboration), and the groups (i.e., teams). For the technology concept, we included several synonyms and related terms to maximize coverage, namely “Social Virtual Reality,” “Social VR,” “Multi-user Virtual Reality,” and “Collaborative Virtual Environments.” We intentionally did not include the term “Virtual Reality” alone, as our focus was on Social VR environments specifically designed to enable interaction and collaboration rather than VR more broadly. For the collaboration concept, we chose a set of keywords focused on people working together: “Collaborative,” “Collaboration,” and “Cooperation.” We chose to add the term “cooperation” alongside “collaboration,” as the two are closely related and often used interchangeably in the literature (Stout and Keast, 2021). Lastly, for the groups concept, we included the words referring to teams: “Team,” “Teams,” “Teamwork,” “Group,” and “Groups.” These words defined the importance of studies analyzing multiple individuals or users within a unit. We combined the synonyms within each concept using the Boolean OR operator and then linked the three main concepts with the AND operator. As a result, our final query was: (“Social Virtual Reality” OR “Social VR” OR “Multi-user Virtual Reality” OR “Collaborative Virtual Environments”) AND (Team OR Teams OR Teamwork OR Group OR Groups) AND (Collaborative OR Collaboration OR Cooperation).
We conducted the first step, “Identification,” using the following three databases: IEEE Xplore Digital Library, ACM Digital Library, and Web of Science. We chose these datasets since they cover the leading conferences and journals on human-computer interaction and team science. We searched for articles in these databases based on their titles, abstracts, and full text. We conducted the final search on February 19, 2025. We included articles published between 2004 and 2024, covering 20 years of research. We limited the included papers to the last 20 years to ensure that our corpus captured the most recent developments, which led to the significant boom in HMD technologies starting around the 2010s, including HTC Vive and Oculus Rift. We also checked earlier studies, which began as early as 1993 in our search scope, and primarily focused on collaboration in virtual environments using screen-based technologies rather than HMDs. For the few early papers involving VR, their contributions were either revisited in later research or shaped by the technological constraints of the time, which limited their applicability to the current landscape of Social VR research. We obtained 892 papers from the IEEE Xplore Digital Library, 972 papers from the ACM Digital Library, and 182 papers from the Web of Science database, resulting in a total of 2,046 papers. Since the ACM Digital Library did not allow users to export the results from their website, we manually copied and pasted the ACM Digital Library’s results to a Google Spreadsheet. We exported the results of all three databases from their websites to a CSV file. We then compiled all the results into a single spreadsheet for analysis. In cases where we found two versions of an article (e.g., a conference paper later extended into a journal article), we only included the most recent version. Based on the titles, we removed 11 duplicate articles. This left us with 2,035 articles for screening. Fig. 2 depicts the trend of paper counts for each year in our search corpus from 2004 to 2024.
Figure 2. The number of search results of the final query in three databases between 2004-2024.
3.2. Eligibility Criteria
Inclusion
We included articles that specifically addressed collaboration within Social VR, focusing on how it can support people and organizations in collaborating and accomplishing work-related tasks. Our inclusion criteria were:
  • Articles must involve two or more human users collaborating in Social VR. Collaboration involves communication, coordination, and sharing of knowledge and resources among participants to achieve a joint objective.
  • Articles must focus on Social VR applications experienced through head-mounted displays (HMDs).
  • The collaborative activity or task described in the paper must have a clear goal relevant to a work setting, such as problem-solving, decision-making, or a creative project. Simply meeting, learning, chatting without a shared work-related purpose, or casual socializing were not considered collaborative tasks for this review.
  • We included only full papers published in journals or conference proceedings to focus on more substantial and in-depth contributions to the literature on collaboration in Social VR.
Exclusion
We decided to exclude articles that were not focused on our research questions or were not accessible to the research team. Our exclusion criteria were:
  • Articles that examined interactions only between virtual agents without any human users. Our interest was in understanding collaboration between real human users.
  • Articles where the VR environment included only a single human participant interacting with virtual agents were also excluded, as this does not constitute a human team.
  • Articles that only examined desktop-based virtual environments or other screen setups.
  • Articles that primarily focused on other forms of extended reality (XR) technologies besides VR, such as augmented reality (AR) and mixed reality (MR). While these XR technologies are related to VR and can also support remote collaboration, they involve different technical setups, interaction modalities, and user experiences. To maintain a clear scope, we limited our review to papers that specifically examined collaboration within fully immersive virtual reality environments using head-mounted displays.
  • Articles primarily focused on (i) educational settings, such as teaching in classrooms or to children; (ii) training either for medical purposes or for specialized fields; and (iii) gaming, whether for entertainment or training purposes. We excluded these papers to concentrate the review on work contexts.
  • Workshop papers, magazines, posters, extended abstracts, and thesis papers. While potentially insightful, these formats were often too brief to provide a substantial contribution.
  • Articles written in languages other than English were excluded to avoid potential misinterpretation.
  • Articles that could not be accessed or downloaded from their original published source.
3.3. Article Selection
The first author of the paper (i.e., the coder) manually screened the titles and abstracts of all 2,035 retrieved articles. The coder proceeded to screen articles for inclusion through a three-stage process. First, the coder reviewed the retrieved articles’ titles and keywords (i.e., level-one screening). The coder retained articles whose titles or keywords met the inclusion and exclusion eligibility criteria, resulting in 1,365 articles excluded and 670 articles selected for the next stage. Then, in the level-two screening, the coder performed a second review that included the articles’ titles, keywords, and abstracts. Based on the inclusion and exclusion criteria, the coder excluded 465 articles and selected 204 articles to read in full. In the third stage (“Eligibility”), the coder read and reviewed each article’s main content in full, excluding 142 articles and selecting 62 articles to build the final corpus. This selection process is explained in more detail in the results section and illustrated in Fig. 3. The coder then presented the selected articles to the other co-authors, who reviewed the coder’s classification in a second cycle together. All authors agreed to analyze the 62 articles, and the coder continued with the data extraction and synthesis stage. Fig. 4 shows the yearly trend of the 62 articles included in our search corpus. The final list of included articles is available in the Supplementary Materials.
Figure 3. PRISMA Flow Diagram of this study. It presents the details of the article selection process.
[…]
PRISMA Flow Diagram of this study. It presents the details of the article selection process.
3.4. Data Extraction and Synthesis
Once the final corpus was established, the coder extracted data to synthesize their characteristics, methods, and findings. Using a Google Spreadsheet, the coder extracted data pertaining to:
  • Year of publication
  • Research type: Bibliographic, Descriptive, Case Study, Laboratory Experiment, Explanatory, or Non-empirical.
  • Evaluation Methodology: Quantitative, Qualitative, or Mixed Methods.
  • VR device type if used in the study (i.e., HTC Vive, Oculus Rifts, Meta Quest).
  • Number of participants recruited in the study.
  • Number of participants per group.
  • Type of experimental task of the study.
The coder also used a Google Doc to take open notes and memos that synthesized the key findings of each article. On average, the coder took 15 notes per article. These notes provided more context and details on the technologies, design processes, and results described in each article. These notes are also available in the Supplementary Materials.
Figure 4. The number of included articles in three databases between 2004-2024.
[…]
The number of included articles in three databases between 2004-2024.
3.5. Thematic Analysis
We conducted a thematic analysis to identify cohesive themes that captured patterns across the final corpus (Braun and Clarke, 2006). The goal of this analysis was to inductively derive themes grounded in the data while minimizing confirmation bias. The coder re-reviewed the articles, including their extracted data and the notes summarizing the main contributions, to generate initial codes. These codes were developed inductively, without reference to pre-existing coding frameworks or theoretical models (Nowell et al., 2017; Braun and Clarke, 2006). The codes were not mutually exclusive since the articles could address multiple codes. The coder iterated the initial codes through multiple rounds of analysis until no new ideas emerged from them. Following this, the refined codes were grouped into themes, which were further revised and renamed over several rounds to improve their clarity and alignment with the corpus data. To enhance reliability and interpretive consensus, all the authors reviewed and discussed the themes and their boundaries in multiple meetings. Ultimately, the final themes were organized into three overarching categories and nine subcategories that captured the breadth and nuance of how Social VR has been studied in the context of collaboration.
4. Results
After applying the methodology described above, 62 of the 2,035 papers were included in our final corpus. Fig.  3 shows the filtering process of the papers in IEEE Xplore Library, ACM Digital Library, and Web of Science databases. The most common reason for excluding articles during the eligibility stage was that they were not related to collaboration in VR—meaning the study either focused on collaboration without using VR (e.g., in 2D platforms), used VR without addressing collaboration, or addressed neither. This exclusion category is labeled in Fig.  3 as ”not related to VR, collaboration, or both” (n = 356). Following this process, Fig. 4 depicts the trend in the number of papers for each year among the 62 included articles from 2004 to 2024.
4.1. Description of the Included Articles
Table 1 provides an overview of research and experimental characteristics of the 62 included articles. Table 1a summarizes the research types, evaluation methodologies, and analysis methods. The majority were laboratory experiments (69.35%), followed by non-empirical papers (12.90%), survey papers (8.07%), and qualitative descriptive studies (4.84%). Three articles (4.84%) fell under other categories, including one case study, one conceptual framework, and one system architecture description.
Regarding evaluation methodology, 33.87% of the studies employed mixed methods, 32.26% used quantitative methods, and 17.74% used qualitative approaches. In 16.13% of the articles, an evaluation methodology was not employed or specified. Among the qualitative methods articles, thematic analysis was the most common method (19.36%), followed by free-text responses (4.84%) and participant experience summaries (3.23%). A subset of studies employed methods such as open coding, observational techniques, or participant experience summaries. On the quantitative side, ANOVA was the most frequently used analysis (19.28%), followed by descriptive statistics (15.67%), t-tests and Wilcoxon tests (12.05% each), and post-hoc or Mann-Whitney U tests (combined 13.24%). Since many studies applied more than one technique, the total number of quantitative methods exceeds the number of papers.
Table 1b summarizes technical and task-related details for the 43 articles that employed laboratory experiments. The most frequently head-mounted display used was the HTC Vive (20.93%), followed by Oculus Quest 2 (16.29%), HTC Vive Pro and HTC Vive Pro Eye (9.30% each), and Oculus Rift (6.99%). Other headsets, including Meta Quest 3, Oculus Go, and early Oculus Quest models, accounted for smaller proportions.
In terms of types of tasks, problem-solving was most common task employed in these articles (23.26%), followed by navigation/ wayfinding (16.28%) and design/construction (16.28%). Other tasks included negotiation/planning (11.62%), meetings/discussions and data analysis (9.30% each), scientific tasks (6.98%), and one workplace simulation (2.33%). Two studies (4.65%) included task types that did not fall under the primary categories outlined in the table.
Lastly, most articles with laboratory studies focused on small teams, including dyads (two-person groups, 79.08%) and triads (16.28%). Only two studies involved more participants per group, one study with four participants, and another with seven participants.
Characteristic Count (%)
Research Type
Laboratory Experiment 43 69.35
Non-empirical 8 12.90
Survey Paper 5 8.07
Descriptive (Qualitative) 3 4.84
Other 3 4.84
Evaluation Methodology
Mixed Methods 21 33.87
Quantitative 20 32.26
Qualitative 11 17.74
N/A 10 16.13
Qualitative Analysis Method
Thematic Analysis 12 19.36
Free-text Response Analysis 3 4.84
Participant Experience Summary 2 3.23
Open Coding Analysis 1 1.61
Observational Evaluation 1 1.61
N/A 43 69.35
Quantitative Analysis Method
ANOVA 16 19.28
Descriptive Statistics 13 15.67
T-test 10 12.05
Wilcoxon Test 10 12.05
Post-hoc Tests 6 7.22
Mann-Whitney U Test 5 6.02
Others 23 27.71
a. Research Types and Analysis Methods (All 62 Studies)
Characteristic Count (%)
VR Device Type
HTC Vive 9 20.93
HTC Vive Pro 4 9.30
HTC Vive Pro Eye 4 9.30
Oculus Rift 3 6.99
Oculus Go 1 2.33
Oculus Quest 2 4.65
Oculus Quest 2 7 16.29
Meta Quest 3 1 2.33
Meta Quest Pro 1 2.33
Others 11 25.58
Experimental Task Type
Problem Solving 10 23.26
Navigation/Wayfinding 7 16.28
Design/Construction 7 16.28
Data Analysis/Visualization 4 9.30
Negotiation/Planning 5 11.62
Meeting/Discussion 4 9.30
Scientific 3 6.98
Workplace Simulation 1 2.33
Other 2 4.65
Participants per Group
Two 34 79.08
Three 7 16.28
Four 1 2.32
Seven 1 2.32
b. Devices, Tasks, and Group Sizes (43 Lab Studies)
Table 1. Overview of research and experimental characteristics of the 62 included articles. Table 1a summarizes the research types, evaluation methodologies, and analysis methods. Table 1b presents VR device types, experimental tasks, and group sizes from 43 lab studies.
4.2. Thematic Analysis
Our thematic analysis revealed three main themes: (1) “Individual Perception and Experience of Collaboration in Social VR,” (2) “Team Dynamics and Collaboration in Virtual Environments,” and (3) “Enabling Collaboration through Social VR Affordances” These themes reflect an increasing focus on employing Social VR to support collaboration by leveraging user experience, team dynamics, and system affordances. Moreover, they uncover significant gaps and future research directions, especially in understanding how specific user interactions and technical elements could effectively enable collaboration in Social VR.
The first theme includes articles addressing the foundational aspects of individual perception and experience within Social VR environments. This theme includes two sub-themes: “Am I Really There?,” which highlights how Social VR can simulate users’ physical presence, enhancing their interaction and engagement with the collaborative task. The second sub-theme, “How Am I Perceived Here?,” explores avatar design and its impact on collaborative dynamics and engagement, emphasizing the importance of avatar customization to shape user interactions and self-expression within virtual spaces.
The second theme focuses on team dynamics and collaboration in immersive virtual environments, examining how Social VR can enhance team collaboration through social presence. This theme includes four sub-themes: “Are We Really Together?,” which explores how social presence contributes to an improved team collaboration; “Is This Collaboration Real?,” which addresses how user perceive the authenticity of collaboration in VR; “Should I Trust this Person?,” which explores trust-building as a critical element of effective virtual collaboration; and “Are We Fully Engaged?,” which considers the role of social engagement in VR collaboration and the different factors that shape user involvement.
The final theme examines how collaboration is supported by the unique affordances of Social VR. These affordances refer to the technical elements and interactions designed to enable realistic and functional 3D virtual environments. This theme includes three sub-themes: “Creating Realistic and Functional Virtual Environments,” which addresses how the design of virtual spaces that mimic real-world settings can improve user experience; “Facilitating Interaction through Virtual Tools and Objects,” which emphasizes the importance of object manipulation and tool use in collaborative tasks; and “Enhancing Immersion through Spatial Audio and Movement,” which underscores how spatial audio and coordinated movement contribute to a more immersive and effective collaborative experience.
A summary of these themes, sub-themes, and their respective articles can be found in Table 2.
Theme N Articles
Individual Perception and Experience of Collaboration in Social VR.
“Am I Really There?” Sense of Presence as a Foundation for Immersive Collaboration. 4 (Oh et al., 2018), (Olaosebikan et al., 2022), (Khojasteh and Won, 2021), (Mei et al., 2021)
“How Am I Perceived Here?” Avatar Design Shaping Collaborative Dynamics and Engagement. 7 (Osborne et al., 2023), (Torro and Pirkkalainen, 2023), (Yang et al., 2024), (Yassien et al., 2021), (Freeman et al., 2020), (Kimmel et al., 2024), (Kocur et al., 2020)
Team Dynamics and Collaboration in Virtual Environments.
“Are We Really Together?” Social Presence Enhancing Team Collaboration. 7 (Olaosebikan et al., 2022), (Dey et al., 2024), (Darbar et al., 2024), (Bovo et al., 2022), (Ide et al., 2020), (Adkins et al., 2024), (Mei et al., 2021)
“Is This Collaboration Real?” Perceiving Realness in Virtual Teamwork. 6 (Torro and Pirkkalainen, 2023), (Maloney and Freeman, 2020), (Osborne et al., 2023), (Bailenson and Yee, 2006), (Park et al., 2024), (Dey et al., 2024)
“Should I Trust this Person?” Enabling Trust as the Glue that Binds Collaborators. 2 (Torro and Pirkkalainen, 2023), (Yassien et al., 2020)
“Are We Fully Engaged?” The Importance of Social Engagement in VR Collaboration. 3 (Xu et al., 2023), (Le Tarnec et al., 2024), (Moharana et al., 2023)
Enabling Collaboration through Social VR Affordances.
Creating Realistic and Functional Virtual Environments. 6 (Mei et al., 2021), (Osborne et al., 2023), (Men and Bryan-Kinns, 2019), (Freiwald et al., 2020), (Weissker and Froehlich, 2021), (Langa et al., 2022)
Facilitating Interaction through Virtual Tools and Objects. 4 (García et al., 2008), (Dey et al., 2024), (Osborne et al., 2023), (Roberts et al., 2005)
Enhancing Immersion through Spatial Audio and Movement. 6 (Weissker et al., 2019), (Immohr et al., 2024), (Ríos et al., 2018), (Immohr et al., 2023), (Weissker, 2020), (Naef et al., 2004)
Table 2. Themes and sub-themes identified through thematic analysis of Social VR research on collaboration. Each row presents a sub-theme, the number of articles (N) associated with it, and its references.
[…]
Themes and sub-themes identified through thematic analysis of Social VR research on collaboration. Each row presents a subtheme, the number of articles (N) associated with it, and representative references. The themes are grouped into three overarching categories reflecting different levels of collaborative experience in Social VR: individual perception, team dynamics, and technological affordances.
4.2.1. Individual Perception and Experience of Collaboration in Social VR
“Am I Really There?” Sense of Presence as a Foundation for Immersive Collaboration.
This sub-theme focuses on individual users’ sense of presence—the subjective feeling of “being there” within a 3D virtual environment (Triberti et al., 2025)—as a foundational condition for collaboration in Social VR. Distinct from social presence, which refers to the perception of others in the environment, sense of presence emphasizes the users’ personal immersion in the virtual space. This experience is often shaped by how intuitively users’ actions and intentions are translated into the virtual environment (Oh et al., 2018; Olaosebikan et al., 2022).
Several studies in our corpus indicated that Social VR can enhance a strong sense of presence by allowing users to act and interact in more natural and embodied ways (Olaosebikan et al., 2022; Khojasteh and Won, 2021; Mei et al., 2021). For instance, Olaosebikan et al. (Olaosebikan et al., 2022) found that scientists collaborating in VR reported a greater sense of immersion and presence compared to those using traditional video conferencing platforms. Participants described their VR experiences as attending a live presentation, highlighting the engaging and spatial qualities of immersive environments (Olaosebikan et al., 2022).
Similarly, (Khojasteh and Won, 2021) conducted a five-session longitudinal study in which users repeatedly collaborated in VR with different partners. Over time, participants reported increased feelings of presence, suggesting that growing familiarity with the virtual environment, task, and interaction patterns contributes to a more immersive experience (Khojasteh and Won, 2021).
Mei et al. (Mei et al., 2021) explored how a co-design activity in VR supported ideation and task engagement. Participants reported that being able to spatially manipulate and visualize design elements within a shared 3D virtual environment enhanced their sense of “being in the task.” Their findings highlight how interactive design tools can strengthen users’ presence through embodied interaction.
In sum, these studies suggest that fostering a strong sense of presence is necessary for effective collaboration in Social VR. When the applications support natural movement, spatial awareness, and meaningful shared activities, users are more likely to feel grounded in the virtual environment and interact intuitively with others. Over time, repeated exposure and growing familiarity with the virtual space can deepen users’ immersion, enabling collaborators to engage more fluidly and confidently.
“How Am I Perceived Here?” Avatar Design Shaping Collaborative Dynamics and Engagement.
This sub-theme examines how avatar design influences users’ experiences in Social VR. We found that multiple articles focused on both how users perceive their own virtual bodies (i.e., self-perception) and how they believe others perceive them. Their main findings show that avatar embodiment plays a key role in how users express themselves, interpret social cues, and participate in collaborative settings. The visual and behavioral fidelity of avatars impacts users’ sense of identity and presence, influencing their confidence, expressiveness, and overall engagement during collaboration.
The articles in this sub-theme demonstrated that avatar characteristics— such as gender, realism, level of abstraction (i.e., how simple or complex the avatar appears), and humanoid forms—significantly shape users’ experiences (Osborne et al., 2023; Torro and Pirkkalainen, 2023; Yang et al., 2024). Gender representation emerged as a recurrent aspect across these studies. For instance, Yassien et al. (Yassien et al., 2021) found that having all the group members’ avatars with the same gender (e.g., all-male or all-female) improved collaboration quality and mutual support, even when users’ real-world gender was different. Similarly, Yang et al. (Yang et al., 2024) demonstrated that users represented by non-human avatars (e.g., fox, robot, duck) reported greater comfort and equal participation than when using communication formats where their real-world appearance was visible, such as face-to-face or video conferencing settings, suggesting that abstract representations may alleviate social pressure and enable more balanced collaboration.
Other studies emphasized the importance of avatar customization and stylization when collaborating in Social VR (Osborne et al., 2023; Freeman et al., 2020). Osborne et al. (Osborne et al., 2023) showed that allowing users to personalize their avatars’ appearance enhanced engagement with the rest of the group. Similarly, Freeman et al. (Freeman et al., 2020), in a descriptive study involving 30 semi-structured interviews (Freeman et al., 2020), demonstrated that enabling self-representation in Social VR supports identity exploration, particularly for users navigating gender or sexual identity. Participants often described their avatars as extensions of themselves, using them to explore self-image in ways that felt both safe and empowering (Freeman et al., 2020).
However, some articles also showed that increased realism in avatar design can complicate users’ sense of self-identification by creating a mismatch between the avatar’s appearance and the user’s internal self-image. Two studies addressed how the “uncanny valley” effect—which occurs when avatars look nearly human but still exhibit subtle imperfections in appearance or movement—made participants feel unnatural or unsettled during the collaboration experience (Osborne et al., 2023; Torro and Pirkkalainen, 2023). Osborne et al. (Osborne et al., 2023) and Torro et al. (Torro and Pirkkalainen, 2023) reported that these unsettling avatars disrupted users’ immersion and hindered their natural interactions. While humanoid avatars can promote familiarity and trust, these findings underscore the importance of addressing the uncanny valley when designing avatars for collaborative experiences in Social VR.
Several articles examined how the fidelity of avatars’ motion and expression influences users’ perceptions of being accurately represented in their social interactions. Kimmel et al. (Kimmel et al., 2024) and Yassien et al. (Yassien et al., 2020) found that full-body avatars with realistic tracking and expressive features (e.g., gaze direction, facial movement) helped users feel more accurately represented. Similarly, Kocur et al. (Kocur et al., 2020) further examined how users believed they were perceived by their teammates in VR. They found that mismatches between how individuals wanted to present themselves and how they believed others perceived them—often due to avatar limitations—undermined users’ confidence and reduced their engagement in group collaboration.
In summary, these articles demonstrate that avatars are not merely cosmetic features but critical components of users’ collaborative experiences in VR. They influence how individuals perceive themselves, their social roles, expressions, and feelings of being acknowledged within shared spaces. As the articles in this sub-theme illustrate, design choices regarding avatar realism, abstraction, gender, and expressiveness can deeply affect users’ comfort, sense of inclusion, and confidence during collaborative tasks in VR.
4.2.2. Team Dynamics and Collaboration in Virtual Environments
“Are We Really Together?” Social Presence Enhancing Team Collaboration.
This sub-theme explores how social presence—the sense of “being with others” in a shared virtual environment (Oh et al., 2018)—facilitates effective team collaboration in Social VR. While the sense of presence refers to a user’s perception of immersion, social presence refers to the awareness of and connection to teammates. It addresses the feeling that others are truly co-located and responsive within the virtual space.
Several articles emphasized that social presence is a foundational social-cognitive state that enables key collaborative behaviors such as turn-taking, mutual coordination, and shared understanding (Olaosebikan et al., 2022; Dey et al., 2024). In a systematic review of Social VR teamwork studies, Dey et al. (Dey et al., 2024) found that social presence strongly influences team performance, citing increased mutual awareness, better turn coordination, and stronger engagement as key team outcomes.
A key component of social presence in VR is mutual attention, which is the ability to perceive others’ attention and actions in real-time (Darbar et al., 2024; Bovo et al., 2022; Ide et al., 2020). For example, Darbar et al. (Darbar et al., 2024) introduced ‘GazeMolVR,’ a system for molecular visualization in collaborative VR that incorporated shared eye-gaze cues. Their findings showed that real-time gaze sharing improved the feeling of connectedness and eased conversational flow, particularly in distributed teams. When collaborators could see where others were looking, they adjusted their own attention more fluidly, making the interaction feel more natural and co-present (Darbar et al., 2024).
The articles also show that social presence is strongly connected to nonverbal awareness and emotional connection. When avatars or systems convey cues such as gaze direction, pointing, or simple gestures, teammates can more easily interpret each other’s intentions (Bovo et al., 2022; Ide et al., 2020). For example, studies such as Bovo et al. (Bovo et al., 2022) and Ide et al. (Ide et al., 2020) implemented visual aids, including cones of vision and symbolic gestures, to help users perceive their collaborators’ focus and actions. These features not only added expressiveness but also enhanced social presence by increasing mutual visibility and acknowledgment, helping users feel noticed and understood by others.
One article highlighted that social presence enhances coordination during multi-party interactions (Adkins et al., 2024). In their evaluation of audio and interaction modalities, Adkins et al. (Adkins et al., 2024) found that teams using spatial audio, compared to non-spatial conditions, experienced smoother turn-taking and reported feeling more “in sync” with one another. These findings suggest that social presence is shaped not only by visual or avatar-based cues but also by auditory and behavioral alignment that supports team coordination.
One last article noted that weak or underdeveloped social presence negatively affects collaboration. In their evaluation of a co-design task, Mei et al. (Mei et al., 2021) found that the absence of avatar reactivity and expressive behavior made interactions feel more like a phone call rather than a shared, embodied collaboration. This diminished sense of social presence reduced user engagement and made communication feel less natural, highlighting the importance of social presence for enabling teamwork.
In summary, these articles demonstrate that social presence serves as the social glue that makes collaboration in Social VR feel natural and connected. The articles show how it fosters mutual attention, real-time responsiveness, emotional connection, and spatial alignment, all of which contribute to a strong sense of togetherness. Ultimately, social presence enables team members to perceive and respond to one another’s actions in real time, supporting cohesive and authentically shared collaboration in VR.
“Is This Collaboration Real?” Perceiving Teamwork’ Realness in Social VR
This sub-theme includes articles that examine how users evaluate the authenticity, seriousness, and naturalness of their collaborative experiences in Social VR. The focus is on whether immersive VR environments support interactions that are not only technically possible but also feel socially valid, natural, and meaningful. Beyond the technical functionality, these articles highlight that users carry expectations shaped by physical-world norms, emotional cues, and social dynamics, which influence how “real” collaboration feels in Social VR (Torro and Pirkkalainen, 2023; Maloney and Freeman, 2020; Osborne et al., 2023).
A key factor in how users evaluate the authenticity of virtual collaboration is whether their partners appear as genuine social actors (Torro and Pirkkalainen, 2023). Torro et al. (Torro and Pirkkalainen, 2023) introduced the concept of the “social presence illusion” to describe how lifelike and responsive avatars make interactions feel more socially authentic. VR systems that support behavioral realism—such as timely gaze, natural gestures, and coordinated responses—enable familiar social rituals (e.g., turn-taking, nonverbal responses) that help elevate collaboration from a purely functional exchange into an interaction that feels intuitively real (Torro and Pirkkalainen, 2023; Bailenson and Yee, 2006).
Building on this, Maloney et al. (Maloney and Freeman, 2020) interviewed long-term Social VR users to understand what made virtual interactions feel meaningful. Participants reported that their collaborations felt more authentic when gestures, feedback, and shared activities aligned with everyday social norms. Many of them expressed frustration with unrealistic avatar features—such as exaggerated cartoon-like hands or faces that could not show realistic emotions—which made the interactions feel artificial and less socially authentic. What users desired was not just visual realism, but interaction fidelity, which translated into responsiveness and expressiveness that matched the tone and rhythm of real-world collaboration.
Other articles in this sub-theme also indicated that misaligned avatar behavior can disrupt the perceived authenticity of collaboration. For example, Osborne et al. (Osborne et al., 2023) noted that technical glitches, such as avatar duplication or unnatural expressions, made interactions feel artificial and awkward. In contrast, when avatars moved and responded in socially expected ways, participants perceived the collaboration as smooth and believable, even when the avatars were abstract in design (Osborne et al., 2023).
Beyond appearance, one article explored how subtle behavioral manipulations could enhance the cohesion and authenticity of collaboration in Social VR (Bailenson and Yee, 2006). In their foundational work, Bailenson et al. (Bailenson and Yee, 2006) introduced the concept of “Transformed Social Interaction” (TSI), demonstrating that altering avatar behaviors (e.g., gaze direction, synchronized gestures) can increase perceived team unity. Their findings suggest that collaborative realism depends not only on shared goals or tasks established in the virtual environment but also on how interaction feels on a social level. By aligning nonverbal cues with social expectations, Social VR applications can strengthen the psychological sense of working as a single team (Bailenson and Yee, 2006).
One study examined how to reinforce the role of realism in a Social VR workplace by eliminating fantastical elements. Park et al. (Park et al., 2024) removed the teleportation feature and required users to navigate through halls and elevators, mirroring how people would walk in a physical office. While this approach increased social encounters and enhanced users’ social presence, it also introduced usability challenges, revealing a tension between realism and practicality (Park et al., 2024). The findings of this article echo broader concerns about how much realism is necessary to foster authentic collaboration without compromising the unique affordances of VR.
Lastly, one article (Dey et al., 2024) showed that Social VR who first engaged in an emotionally intense solo experience approached a subsequent group task with more seriousness and coordination. The realism of the initial simulation made the following collaboration feel more consequential and authentic. Dey et al.’s findings (Dey et al., 2024) suggest that initial immersive experiences can reshape how users’ perceptions of group work in Social VR, making virtual collaboration feel more meaningful, important, and socially grounded.
In summary, these articles demonstrate that the authenticity of collaboration in Social VR is not just a result of visual fidelity or technical features, but on how well interactions mirror the rhythms, responsiveness, and emotional tone of real-world teamwork. When Social VR users feel that their actions have an impact and that their collaborators are genuinely engaged, the experience transcends the simulation and becomes a psychologically meaningful, socially credible form of collaboration.
“Should I Trust this Person?” Enabling Trust in Social VR
This sub-theme explores how trust, an essential yet frequently overlooked foundation of effective collaboration, is developed and sustained in Social VR. In co-located settings, trust often emerges through informal interactions and subtle nonverbal cues. However, these trust-building mechanisms are often limited or absent in remote communication settings. Nevertheless, the reviewed articles show that Social VR can foster interpersonal trust through its distinctive affordances, including embodied interaction, avatar behavior, and shared spatial context (Torro and Pirkkalainen, 2023; Yassien et al., 2020).
Torro et al. (Torro and Pirkkalainen, 2023) developed a design theory explaining how Social VR can support trust-building in virtual teams. Grounded in social exchange theory, which emphasizes reciprocal interactions as the foundation of relationships (Cook et al., 2013), they proposed that trust emerges through two key pathways in Social VR: a peripheral route, driven by first impressions (e.g., avatar appearance, social cues), and a central route, based on deeper assessments of a partner’s competence, integrity, and goodwill (Torro and Pirkkalainen, 2023). The article emphasized that immersive Social VR applications designed to support both verbal and nonverbal communication can enhance trust through both routes. For example, avatars that mimic natural gestures, proximity-based conversations, and realistic movement can help promote social bonding. Moreover, features like virtual break rooms with spatial properties can support casual, unstructured interactions, which are often absent in structured video calls or text-based platforms.
Yassien et al. (Yassien et al., 2020) highlighted another key dimension of trust in Social VR: personal space. This article found that users felt safer and more connected when their virtual personal space was respected. Trust was undermined when avatars came too close or moved in ways that felt erratic or unpredictable, causing discomfort and breaking the sense of social ease. The researchers recommended designing Social VR applications with mechanisms to prevent collisions and unwanted proximity between avatars, reinforcing the importance of comfort and perceived safety in virtual interactions.
In summary, these two articles suggest that trust can be supported through both social signaling and environmental design, such as expressive avatars, respectful personal space, and opportunities for informal interaction.
“Are We Fully Engaged?” The Importance of Social Engagement in VR Collaboration.
This sub-theme focuses on social engagement, the extent to which collaborators in Social VR remain emotionally and communicatively connected during a shared task (Xu et al., 2023). Social engagement goes beyond mere attention or performance, it reflects how present, expressive, and emotionally involved individuals feel in relation to their teammates (Xu et al., 2023). The reviewed articles identified design strategies that support behavioral cues of engagement and help sustain interpersonal connections throughout the Social VR collaboration (Le Tarnec et al., 2024; Xu et al., 2023; Moharana et al., 2023).
One strategy identified in this sub-theme involves enhancing avatar behaviors to foster social engagement. Le et al. (Le Tarnec et al., 2024) introduced an augmented avatar system that generated subtle nonverbal cues—such as head nods and posture shifts—even when users did not perform them directly. These enhancements made collaborators appear more responsive and involved, leading to more natural and engaging teamwork (Le Tarnec et al., 2024). Similarly, Xu et al. (Xu et al., 2023) studied older adults co-planning virtual travel experiences using VR and found that social engagement deepened when participants shared personal memories and life stories related to the destinations. These moments of self-disclosure fostered warmth and mutual interest, highlighting that social engagement can be supported not only through interface design but also through content that invites personal connection and emotional expression.
Another article explored physiological synchrony as a deeper measure of engagement. Moharana et al. (Moharana et al., 2023) tracked heart rate patterns during a collaborative memory game and found that teams with higher physiological synchrony reported greater feelings of connection, presence, and team flow. These findings suggest that social engagement may operate at both cognitive and emotional levels and that VR environments can help capture or even facilitate these more subtle social dynamics.
Taken together, these articles show that social engagement in VR collaboration extends beyond visual realism or task efficiency. It hinges on participants feeling seen, acknowledged, and emotionally connected, which are qualities that can be nurtured through expressive avatars, thoughtful interaction design, and moments that invite genuine interpersonal connections. Therefore, designing for sustained social engagement is essential for effective, human-centered teamwork in Social VR applications.
4.2.3. Enabling Collaboration through Social VR Affordances
Creating Realistic and Functional Virtual Environments.
This sub-theme explores how the spatial layout, visual realism, and responsiveness of the virtual environment influence users’ interactions and support collaboration in Social VR. Rather than focusing on tools or input mechanisms, the reviewed articles emphasize how the environment’s visual design, structure, and adaptability influence where and how users interact. The findings suggest that effective virtual environments in Social VR balance familiar spatial cues with flexible design elements, enabling collaborative behaviors that can surpass those possible in the physical world (Mei et al., 2021; Osborne et al., 2023; Men and Bryan-Kinns, 2019; Freiwald et al., 2020; Weissker and Froehlich, 2021; Langa et al., 2022).
Two articles found that environmental realism is most valued when it reinforces familiar spatial dynamics. Mei et al. (Mei et al., 2021) found that the absence of environmental sound, realistic tools, or background detail made the virtual workspace feel less functional to the users. However, the articles discussed that greater realism does not necessarily lead to a better user experience. Osborne et al. (Osborne et al., 2023) proposed a typology of Social VR environments to support users’ engagement: Skeuomorphic, environments designed as realistic office-like spaces; Experimental, environments designed as creative and open-ended spaces, and Prefab-based, which are template-driven but customizable by the users. Their findings suggested that adaptability is often more effective than fidelity. Therefore, environments in Social VR should be tailored to users’ goals and team dynamics to better support their collaborations, instead of focusing solely on visual realism.
One article explains how layout and spatial configuration shape users’ access, privacy, and coordination in Social VR. In a study on collaborative music creation, Men et al. (Men and Bryan-Kinns, 2019) found that participants favored publicly visible personal spaces that balanced individual work zones with group awareness. They also revealed trade-offs in spatial arrangements: positioning avatars side-by-side offered shared perspectives but limited equal access to shared interactive virtual tools, while face-to-face setups improved visibility for all participants but required redesigned interfaces to maintain usability. These findings underscore how spatial affordances—such as visibility, proximity, and orientation—can directly impact collaborative dynamics in Social VR applications.
Two articles highlighted how social VR’s affordances—such as spatial layout and 3D visibility cues—support team coordination by aligning users’ viewpoints and actions (Freiwald et al., 2020). Freiwald et al. (Freiwald et al., 2020) compared three visual indicators for representing users’ viewpoints during a collaborative task. They found that displaying a live feed of each user’s view reduced their task’s error rates, while adding a 3D cone to indicate the boundaries of each user’s viewpoint enhanced their social presence. Similarly, Weissker et al. (Weissker and Froehlich, 2021) developed a group navigation technique that allowed collocated users to teleport together while preserving their relative positions. This feature helped teams stay spatially coordinated as they explored larger virtual environments (Weissker and Froehlich, 2021).
Lastly, one study addressed the technical setup required to scale larger groups and collaborative spaces in Social VR. Langa et al. (Langa et al., 2022) developed a Social VR application for multiparty volumetric meetings, focusing on accessibility and stability for larger groups. While their evaluation involved groups of two and four, the system was designed to scale beyond small teams without sacrificing responsiveness or presence. Their use of volumetric user representations—captured and rendered in 3D—enabled more lifelike, embodied interactions than traditional avatars, offering a promising direction for realistic and scalable Social VR collaboration (Langa et al., 2022).
In summary, these articles suggest that the most effective Social VR applications do not simply mimic the real world. Instead, they introduce new affordances that are not possible to afford in face-to-face meetings—such as view-sharing, spatial partitioning, and adaptable layouts—to support effective collaboration. Employing these immersive affordances to augment team members’ abilities reflects a core principle in CSCW: designing systems for beyond being there by leveraging the unique possibilities of virtual environments rather than constraining them to real-world metaphors.
Facilitating Interaction through Virtual Tools and Objects.
Articles in this sub-theme examine how virtual tools and manipulable objects enable collaboration in Social VR. Unlike traditional online collaboration platforms that rely on screen-based interactions, Social VR allow users to engage with sharedtools using hand movements, gaze direction, and spatial movement, making teamwork feel more tangible and embodied. Across the reviewed articles, tools such as pens, whiteboards, 3D models, and laser pointers were not only instrumental to complete tasks but also served a social function, enabling users to coordinate actions, maintain mutual awareness, and collaboratively build shared outcomes (García et al., 2008; Dey et al., 2024; Osborne et al., 2023).
Two articles examined closely coupled collaborative tasks that required users to coordinate in real-time by synchronizing and sequentially manipulating shared objects (Roberts et al., 2005; García et al., 2008). Roberts et al. (Roberts et al., 2005) designed a Gazebo-building task in which participants had to pass, plan, and co-assemble components in real-time. Garcia et al. (García et al., 2008) addressed the technical challenges in network consistency and inconsistent feedback synchronization across users in VR by introducing visual cues—such as arrows and fall indicators—that helped participants anticipate object behavior and reduced errors during their interactions.
Other articles emphasized the importance of tool accessibility and ease of use for effective collaboration in Social VR. While certain features in VR (e.g., sticky notes, drawing tools) supported brainstorming and shared design tasks, platform-specific constraints (e.g., restricted permissions, complex setup processes) often hindered the collaboration experience (Osborne et al., 2023). Osborne et al. (Osborne et al., 2023) noted that AltspaceVR provided relatively few built-in tools for groups and frequently relied on external software, which sometimes disrupted users’ interaction flow. Complementing these results, Dey et al. (Dey et al., 2024) underscored how direct interaction with virtual objects in Social VR enhances immersion, making the workspace feel more tangible. Their findings suggest that physical engagement with virtual tools not only aids task performance but also strengthens users’ sense of connection to the team and the task (Dey et al., 2024).
Taken together, these articles demonstrate that the design and usability of virtual tools directly affect the collaborative dynamics in Social VR. Intuitive and accessible tools can support effective teamwork by facilitating smoother interactions, idea exchange, and action coordination in coupled tasks. In contrast, complex, restrictive, or poorly designed tools introduce friction to their work and user experiences in VR. Applications that allow multiple users to jointly control, annotate, or manipulate virtual objects foster more balanced participation and fluid idea exchange. Conversely, Social VR applications with limited multi-user access or rigid fixed interaction patterns can reinforce bottlenecks or discourage contribution.
Enhancing Immersion through Spatial Audio and Movement.
The last sub-theme examines how audio and movement technologies in Social VR contribute to a deeper sense of immersion and facilitate user coordination. Affordances such as spatial audio and group locomotion help users feel oriented, responsive, and connected within the shared virtual environment, supporting both awareness and collaboration (Weissker et al., 2019; Immohr et al., 2024; Ríos et al., 2018).
Two articles demonstrated that spatial audio helps anchor users in the virtual environment by simulating the direction and proximity of voices and ambient sounds. Weissker et al. (Weissker et al., 2019) found that using binaural spatial audio—which simulates 3D sound perception by delivering slightly different signals to each ear—improved triadic collaboration by enhancing social presence, turn-taking behavior, and attentional alignment. Participants reported being more aware of their teammates’ positions and smoother conversational flow compared to non-spatial audio conditions (Weissker et al., 2019). However, Immohr et al. (Immohr et al., 2023) found that spatial audio did not significantly improve performance in a negotiation task, suggesting that the effectiveness of audio design depends on the type of task and the dynamics of user interaction. The processing and presentation of spatial audio also influence how immersive and credible virtual interactions feel. Immohr et al. (Immohr et al., 2024) evaluated different binaural rendering methods and found that more realistic spatial cues improved both the perceived authenticity of communication and the naturalness of user responses. These subtle auditory refinements encouraged more lifelike exchanges, particularly in tasks where directional cues and conversational nuance played a critical role.
Movement also plays an important role in collaborative tasks within Social VR, particularly those involving co-navigation or shared spatial exploration. Rios et al. (Ríos et al., 2018) found that synchronizing avatar footstep sounds and animations with users’ real movements enhanced spatial awareness. Participants navigated more accurately and walked closer to objects and collaborators. Although participants’ subjective sense of presence did not change significantly, behavioral indicators suggested stronger alignment between users and their avatars (Ríos et al., 2018).
The reviewed articles introduced a range of group navigation techniques to address the challenges of coordinating movement in collaborative VR applications (Weissker, 2020; Naef et al., 2004; Weissker et al., 2019). For instance, Weissker et al. (Weissker, 2020) developed a group teleportation system that preserved user formations during their jumps. Building on this idea, Naef et al. (Naef et al., 2004) developed the “blue-c API,” which enabled synchronized group movement, allowing users to navigate the virtual environment together without losing cohesion. Weissker et al. (Weissker et al., 2019) developed “Multi-Ray Jumping,” a system designed for collocated groups that offered directional previews, seamless teleportation, and enhanced mutual awareness during shared navigation. These articles show that group movement in VR is not only about getting from one place to another but also about maintaining spatial coordination and shared context to support effective teamwork (Naef et al., 2004; Weissker et al., 2019).
In summary, these articles illustrate how spatial audio and group movement features enhance the authenticity and fluidity of collaboration in Social VR. When thoughtfully designed, these elements can improve how teams communicate, stay together, and navigate virtual spaces using Social VR.
5. Discussion
This review explores two decades of research on Social VR as a medium for supporting collaboration and teamwork. Through a systematic analysis, we characterized how researchers have designed, evaluated, and applied Social VR technologies to enhance remote collaboration. Using thematic analysis, we identified three overarching themes that capture how the literature has conceptualized and evaluated collaboration in social VR applications (see Table 3).
In the following subsections, we discuss the key findings from this review, highlighting how Social VR enables new forms of presence, engagement, and coordination. We reflect on how these findings extend, complement, or challenge prior research in remote collaboration—particularly computer-mediated communication (CMC) and early VR applications—and outline how Social VR introduces new questions and tensions that require more perspectives in CSCW.
Themes Description Key Insights
Individual Perception and Experience
Am I Really There? Examines users’ sense of presence as a foundation for immersive collaboration Presence increases with natural interaction and repeated use. Physical engagement with tasks enhances immersion. Familiarity over time deepens presence.
How Am I Perceived Here? Explores the impact of avatar design on collaborative dynamics Avatar customization supports engagement and identity. Balance needed between realism and avoiding the uncanny valley. Expressive cues shape confidence and inclusion.
Team Dynamics and Collaboration
Are We Really Together? Focuses on how social presence enhances team collaboration Full-body avatars and non-verbal cues increase connectedness. Spatial and emotional alignment improve collaboration. Social presence fosters mutual responsiveness.
Is This Collaboration Real? Addresses the perception of authenticity in virtual teamwork Natural behaviors and responsiveness feel more authentic. Collaboration quality tied to social authenticity. Realism must support, not hinder, collaborative flow.
Should I Trust This Person? Examines trust-building in virtual collaboration Only two studies explored trust directly. Trust built through social cues and personal space. Informal interactions and mimicry enhance trust.
Are We Fully Engaged? Studies the importance of social engagement Non-verbal behaviors support engagement. Shared storytelling and personal content deepen connection. Physiological synchrony linked to team cohesion.
Enabling Collaboration through Social VR Affordances.
Creating Realistic and Functional Virtual Environments Explores the design of functional virtual spaces Environments should go beyond physical replication. Layouts shape access, privacy, and coordination. Customization enhances collaboration effectiveness.
Facilitating Interaction through Virtual Tools and Objects. Examines the object manipulation and tool use Tools support both task performance and coordination. Accessibility and ease of use are critical for collaboration. New tools should support shared, dynamic interaction.
Enhancing Immersion through Spatial Audio and Movement Focuses on spatial audio and movement Spatial audio improves presence and turn-taking. Coordinated movement helps group alignment. Immersive features must match the task context.
Table 3. Overview of Themes and Sub-themes with Descriptions and Key Insights.
5.1. RQ1: Diverse Research from Design to Evaluation in Social VR
The reviewed corpus reveals a growing interest in understanding how Social VR can enhance collaboration. Building on foundational CSCW research that emphasized the importance of conceptualizing distance (Olson and Olson, 2000) and place (Harrison and Dourish, 1996) in virtual communication, the articles examined Social VR as a new medium for group interaction. These articles depart from other major strands of VR research focused on gaming or education. Over the two past decades, advances in hardware, software, and networking have enabled more robust multi-user experiences and triggered new opportunities to create more Social VR experiences. Notably, the rise in research on collaborative Social VR since 2022—influenced by the COVID-19 pandemic and the surge of interest in metaverse technologies—has spurred new systems and theoretical conversations about enabling teamwork in immersive 3D environments.
Our systematic review also highlights how foundational constructs in VR—such as presence, social presence, and the uncanny valley—remain central to evaluating collaboration in Social VR. For example, while the uncanny valley is a known challenge in VR, the reviewed articles show how subtle imperfections in avatar realism can disrupt trust, make users feel socially uneasy, and even affect team cohesion, issues that matter less in entertainment contexts. While these concepts have traditionally been used in HCI and communication studies to assess user experience, their application is expanding to multi-user contexts. As VR becomes more embedded in collaboration, researchers use these frameworks to understand group perception, interaction, and relationships. Bridging these constructs with perspectives from team science (Mathieu et al., 2017) and groups in CSCW (Harris et al., 2019) can offer new theoretical pathways for understanding immersive social dynamics.
One of the key insights is that every design component in Social VR can significantly influence collaboration. Unlike traditional screen-based platforms, Social VR requires full embodied engagement and attachment to the simulated experience. Its immersive nature means that even seemingly minor design decisions—such as avatar posture, hand alignment, spatial audio calibration, or tool responsiveness—can affect users’ comfort, perception, and collaborative performance. While well-designed affordances help users feel grounded and engaged in the virtual workspace, poorly implemented features can lead to disorientation, misalignment, and disengagement.
These challenges are amplified by the immersive nature of VR. While users in video conferencing tools like Zoom can multitask, momentarily disengage, or rearrange their screens with minimal disruption, in Social VR pausing or stepping away from a group session often requires removing the headset or exiting the application, causing a complete break in presence and team cohesion. This absence of pause and exit mechanisms highlights how immersive systems create new forms of “collaborative friction” that do not exist in other communication environments (Shivener and Caravella, ). These “frictions” make Social VR applications more likely to usability breakdowns, reticence, and user frustration. As our review shows, the collaborative experience is shaped by an interplay of multiple factors at the individual, team, and environmental level, including the user’s physical setup, device usability, and familiarity with immersive technologies. Even subtle issues, like delayed hand tracking or disorienting locomotion, can lead to breakdowns in trust, participation, or group cohesion. All of these factors are tightly coupled with the application’s design choices.
Unlike web-based communication platforms—which often do not include the devices as part of their design—building Social VR applications spans multiple technical layers, from HMD ergonomics to affordance design for social interaction. Not having these different layers aligned and appropriately designed can lead to more user frustration. For example, a confusing teleportation system could hinder team navigation, reduce situational awareness, and create participation asymmetries. Earlier collaboration systems often treated social dynamics and interface design as separate concerns, but the articles show that these factors are deeply intertwined. As such, supporting effective collaboration in Social VR requires an integrated approach that considers both technological and social design elements. Developers and designers must ensure that the entire system works cohesively to support seamless and equitable collaboration.
5.2. RQ2: Gaps and future directions
While the reviewed articles show the potential of Social VR for collaboration, they also expose several gaps that should be addressed to unlock its full potential. Most articles focused on short-term sessions or dyads, limiting our understanding of how immersive collaboration unfolds over extended periods or within large, diverse teams. Only a few studies examined repeated use, showing that presence can improve over time (Khojasteh and Won, 2021). Future research should explore long-term use cases and organizational contexts to better assess how Social VR supports ongoing coordination, relationship development, and task management over time.
Another major gap is the lack of standardized frameworks for evaluating collaboration in Social VR. Most studies relied predominantly on self-reported data, with limited behavioral or trace data. While CMC research has long used message logs or performance data, few articles in our review leverage these metrics. More robust, multi-modal methods are needed to capture the complexity of Social VR collaboration. Moreover, most samples were relatively small and homogeneous, limiting generalizability. Future research should examine how user diversity shapes collaboration.
Trust, while fundamental to collaboration, was notably underexplored in these articles. Only two studies in our corpus directly examined how trust develops in Social VR (Torro and Pirkkalainen, 2023; Yassien et al., 2020). These highlighted how avatar expression, proximity, and personal space influence trust. Unlike CMC platforms, where trust often develops via sustained text or asynchronous interactions, trust in VR is embodied and experiential. Understanding how team dynamics unfold in Social VR applications calls for new research on how immersive platforms can contribute to forming trust.
While many studies aimed to recreate physical-world dynamics, many articles encouraged designers to think beyond replication. As illustrated by the articles, Social VR can support forms of collaboration that transcend physical limitations, such as manipulating unreal objects, customizing users’ bodies, and using AI features and tools in immersive ways. Inspired by (Grønbæk et al., 2024; Irlitti et al., 2024), future research should study how immersive collaboration can benefit from VR’s ability to break physical restrictions: supporting asymmetric views, invisible tools, or gravity-defying interactions that foster creativity and engagement. This aligns with the “Beyond Being There” perspective (Hollan and Stornetta, 1992), which advocates for systems that not only simulate co-located interaction but also go further by introducing novel capabilities not possible in physical space.
The articles also raised concerns about usability and fatigue. Many users reported dizziness or discomfort after prolonged VR sessions. Hybrid solutions that combine immersive and non-immersive modes might offer more sustainable approaches to collaboration, especially for long-term or high-frequency use. For instance, teams might brainstorm using VR and then switch to 2D tools for task refinement. Adaptive systems that respond to user context could leverage the benefits of immersion while mitigating its limitations. Furthermore, as Social VR adoption grows, privacy and security concerns must be considered. Sensitive data during collaborations requires mechanisms that protect personal data while preserving presence and immersion.
As Social VR and Mixed Reality evolve, CSCW researchers should move beyond comparing these platforms to traditional online communication tools. Instead, future work should examine how new possibilities provided by immersive systems, such as enabling users to customize shared spaces, collaborating with embodied AI agents (Malik et al., 2020; Wang et al., 2020), and supporting new forms of teamwork that go beyond simulating in-person meetings.
5.3. Limitations and Future Work
It is important to acknowledge the limitations of this paper. First, our systematic review was scoped to three databases only. We did not include articles indexed in other systems (e.g., Scopus, Google Scholar), pre-prints (e.g., arXiv), or non-academic sources. Second, the review focused on papers from the last 20 years. Earlier work could offer additional context on the evolution of Social VR. Third, the inclusion process and thematic analysis were conducted by only one researcher, introducing bias to the selection process. We established meetings to revise and discuss the researcher’s decisions and findings. Lastly, future reviews could explore how other Extended Reality technologies have contributed to collaboration. Despite these limitations, our review advances the field by consolidating emerging work on embodied, immersive collaboration and offering a roadmap for future research in CSCW.
6. Conclusion
This systematic literature review provides a comprehensive analysis of how Social VR has been studied as a medium for collaboration. Drawing from 62 relevant articles, we identified key themes illustrating how Social VR facilitates collaboration at the individual, group, and environmental levels. The findings highlight its potential to enhance teamwork by fostering immersion, enabling expressive and customizable avatars, and facilitating social presence through nonverbal and spatial cues. Yet, important gaps remain in studying long-term collaboration dynamics, team-level processes such as trust-building, and the role of avatar design on shaping self-presentation and group cohesion.
Based on the themes and findings, we encourage researchers and practitioners to investigate sustained collaboration across diverse teams, develop standardized evaluation frameworks, and embrace the unique affordances of Social VR to create more engaging, inclusive, and transformative collaborative experiences.